OpenStudy (anonymous):

Find all roots of the polynomial f(x)= x^4-3x^3+4x^2-6x+4

OpenStudy (lgbasallote):

do you mean +4 at the far end there?

OpenStudy (anonymous):

Yes! oops!

OpenStudy (lgbasallote):

hmm okay..are you familiar with remainder theorem and rational roots theorem? or no?

OpenStudy (anonymous):

No :/

OpenStudy (lgbasallote):

then i guess you have to factor this out the hard way...you know how to do long division right?

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

hmm are you sure this question is right? im not getting a factored form... :/

OpenStudy (anonymous):

Nope I messed up I apologize

OpenStudy (lgbasallote):

lol =))))

OpenStudy (anonymous):

There is the re-re-edited

OpenStudy (lgbasallote):

ahh better...okay...start by dividing it by x-2 and tell me what you get

OpenStudy (anonymous):

I'm reading about the factor therorem would my division look like this \[2 1 1 0 0 \div -2\]

OpenStudy (lgbasallote):

hmm?

OpenStudy (anonymous):

I'm most likely very wrong

OpenStudy (lgbasallote):

where did you get 2110 / -2?

OpenStudy (lgbasallote):

the factor theorem is that if x - 2 is a factor that means f(2) = 0

OpenStudy (anonymous):

I'm trying to look at an example in my book, but I'm failing

OpenStudy (lgbasallote):

here's an example.. x^2 - 2x + 1 i want to know the factors to this using factor theorem i'll try x + 1 so if f(-1) = 0 that means x + 1 is a factor f(-1) = (-1)^2 - 2(-1) + 1 f(-1) = 1 + 2 + 1 f(-1) = 3 it's not 0 so x+ 1 is not a factor. now i'll try x - 1 if x - 1 is a factor f(1) should be 0 f(1) = (1)^2 - 2(1) + 1 f(1) = 1 - 2 + 1 f(1) = 0 that means x - 1 is a factor of x^2 - 2x + 1 make sense?

OpenStudy (anonymous):

Yeah